# Avoiding Catastrophe in Online Learning by Asking for Help

### v3

@@ 3 Model
If $\mu(s_{t},a_{t})\in[0,1]$ is the chance of avoiding catastrophe on time step $t$ (conditioned on no prior catastrophe), then $\prod_{t=1}^{T}\mu(s_{t},a_{t})$ is the agent’s overall chance of avoiding catastrophe. [^7] Conditioning on no prior catastrophe means we do not need to assume that these probabilities are independent (and if catastrophe has already occurred, this time step does not matter). This is due to the chain rule of probability. For a fixed state sequence $\mathbf{s}$ and agent actions $\mathbf{a}=(a_{1},\dots,a_{T})$, the agent’s regret is
$R_{T}(\mathbf{s},\mathbf{a},\mu,\pi^{m})=\prod_{t=1}^{T}\mu^{m}(s_{t})-\prod_{t=1}^{T}\mu(s_{t},a_{t})$
We will usually write $R_{T}=R_{T}(\mathbf{s},\mathbf{a},\mu,\pi^{m})$ for brevity. We will study the expected regret over any randomness in $\mathbf{s}$ and/or $\mathbf{a}$. We desire subconstant worst-case regret: the total (not average) expected regret should go to 0 for any $\mu$ and $\pi^{m}$. Formally, we want $\lim_{T\to\infty}\sup_{\mu,\pi^{m}}\operatorname*{\mathbb{E}}[R_{T}]=0$.
@@
VC dimension [38] and Littlestone dimension [26] are standard measures of learning difficulty which capture the ability of a hypothesis class (in our case, a policy class) to realize arbitrary combinations of labels (in our case, actions). We omit the precise dimensions since we only utilize these concept via existing results. See [34] for a comprehensive overview.
@@
The diameter of $X\subseteq S$ is defined by $\operatorname*{diam}(X)=\max_{x,y\in X}||x-y||$. All logarithms and exponents are base $e$ unless otherwise noted.
@@ 4 Avoiding catastrophe is impossible in general
We begin by showing that in general, any algorithm with sublinear mentor queries has arbitrary poor regret in the worst-case, even when states are i.i.d. on $[0,1]$. The result also holds even if the algorithm knows $L$ and $\mathbf{s}$ ahead of time. We use $\operatorname*{\mathbb{E}}_{\mathbf{s}\sim U^{T},\mathbf{a}}$ denote the expectation over both $\mathbf{s}$ and any additional randomness in the algorithm.
@@
The worst-case expected regret of any algorithm with sublinear queries goes to 1 as $T$ goes to infinity. Formally, $\lim_{T\to\infty}\ \,\sup_{\mu,\pi^{m}}\ \,\operatorname*{\mathbb{E}}_{\mathbf{s}\sim U^{T},\mathbf{a}}\ \,[R_{T}]=1$.
@@ 4.1 Intuition
We partition $S$ into equally-sized sections that are “independent” in the sense that querying a state in section $i$ gives you no information about section $j$. The number of sections is determined by a function $f:\mathbb{N}\to\mathbb{N}$ that we will choose. If $|Q_{T}|\in o(f(T))$, most of these sections will never contain a query. When the agent sees a state in a section not containing a query, it essentially has to guess, meaning it will be wrong a constant fraction of the time. Figure 3 fleshes out this idea.
$s$$\mu$  $\mu(s,0)$$\mu(s,1)$$\cfrac{L}{2f(T)}$$\frac{1}{f(T)}$$\frac{2}{f(T)}$$\frac{3}{f(T)}$$\frac{f(T)-1}{f(T)}$$\dots$1001111
Figure 3: An illustration of the construction we use to prove Theorem 4.1 (not to scale). The horizontal axis indicates the state $s\in[0,1]$ and the vertical axis indicates the payoff $\mu(s,a)\in[0,1]$. The solid line represents $\mu(s,0)$ and the dotted line represents $\mu(s,1)$. In each section, one of the actions has the optimal payoff of 1, and the other action has the worst possible payoff allowed by $L$, reaching a minimum of $1-\frac{L}{2f(T)}$. Crucially, both actions result in a payoff of 1 at the boundaries between sections: this allows us to “reset” for the next section. As a result, we can freely toggle the optimal action for each section independently.
@@
A natural idea is to try $f(T)=T$, but this doesn’t quite work: even if the agent chooses wrong on every time step, the minimum payoff is still at least $1-\frac{L}{2T}$, and $\lim_{T\to\infty}\prod_{t=1}^{T}\left(1-\frac{L}{2T}\right)=\lim_{T\to\infty}\left(1-\frac{L}{2T}\right)^{T}=e^{-L/2}$. In order for the regret to approach 1, we need $f(T)$ to be asymptotically between $|Q_{T}|$ and $T$ (such $f$ must exist since $|Q_{T}|\leq g(T)\in o(T)$). This leads to the following bound: $\prod_{t=1}^{T}\mu(s_{t},a_{t})\leq\big{(}1-\frac{L}{\Theta(f(T))}\big{)}^{\Theta(T)}$. When $f(T)\in o(T)$, the right hand side converges to 0, while $\prod_{t=1}^{T}\mu^{m}(s_{t})=1$. In words, the agent is nearly guaranteed to cause catastrophe, despite the existence of a policy which is guaranteed to avoid catastrophe.
@@ 4.2 Formal definition of construction
Let $S=[0,1]$ and $\mathcal{D}_{t}=U$ for each $t\in[T]$. Assume that $L\leq 1$; this will simplify the math and only makes the problem easier for the agent. We define a family of payoff functions parameterized by a function $f:\mathbb{N}\to\mathbb{N}$ and a bit string $\mathbf{x}=(x_{1},x_{2},\dots,x_{f(T)})\in\{0,1\}^{f(T)}$. The bit $x_{j}$ will denote the optimal action in section $j$. Note that $f(T)\geq 1$ and since we defined $\mathbb{N}$ to exclude 0.
@@
$\mu_{f,\mathbf{x}}(s,a)=\begin{cases}1&\textnormal{ if }a=x_{j(s)}\\ 1-L\left(\cfrac{1}{2f(T)}-|m_{j(s)}-s|\right)&\textnormal{ if }a\neq x_{j(s)}\\ \end{cases}$
Let $\pi^{m}$ be any policy which is optimal for $\mu_{f,\mathbf{x}}$. Note that there is a unique optimal action for each $s_{t}$, since each $s_{t}$ belongs to exactly one $S_{j}$; formally, $\pi^{m}(s_{t})=x_{j(s_{t})}$.
For any $\mathbf{x}\in\{0,1\}^{f(T)}$, $\mu_{f,\mathbf{x}}$ is piecewise linear (trivially) and continuous (because both actions have payoff 1 on the boundary between sections). Since the slope of each piece is in $\{-L,0,L\}$, $\mu_{f,\mathbf{x}}$ is Lipschitz continuous. Thus by Proposition E.1, $\pi^{m}$ satisfies local generalization.
@@ 5 Avoiding catastrophe given finite VC or Littlestone dimension
Theorem 4.1 shows that avoiding catastrophe is impossible in general, which is also true in online learning. What if we restrict ourselves to settings where standard online learning is possible? Specifically, we assume that $\pi^{m}$ belongs to a policy class $\Pi$ where either (1) $\Pi$ has finite VC dimension $d$ and $\mathbf{s}$ is $\sigma$-smooth or (2) $\Pi$ has finite Littlestone dimension $d$ [^8] Recall from Section 1.3 that standard online learning becomes tractable under either of these assumptions.
This section presents a simple algorithm which guarantees subconstant regret and sublinear queries under either of those assumptions. Our algorithm needs to know $\Pi$, as is standard in online learning. The algorithm does not need to know $\sigma$ (in the smooth case) or $L$, and can handle an unbounded state space (the number of queries simply scales with the maximum distance between observed states). For simplicity, we initially prove our result for $A=\{0,1\}$. Appendix C extends our result to many actions using the“one versus rest” reduction. [^9] For each action $a$, we learn a binary classifier which predicts whether $\pi^{m}(s)=a$. If every binary classifier is correct, we can correctly determine $\pi^{m}(s)$. See, e.g., Chapter 29 of [34].
@@ 5.1 Intuition behind the algorithm
Algorithm 1 has two simple components: (1) run a modified version of the Hedge algorithm for online learning, but (2) ask for help in unfamiliar states (specifically, when the current state is far from any queried state with the same action under the proposed policy). Hedge ensures that the number of time steps where the agent’s action doesn’t match the mentor’s is small, and asking for help in unfamiliar states ensures that when we do make a mistake, the cost isn’t too high. This algorithmic structure seems quite natural: mostly follow a baseline strategy, but ask for help when out-of-distribution.
Simple operations.
The algorithm does not require detailed access to the state embedding, instead relying on two simple operations: evaluating a policy on a particular state, and computing a nearest neighbor distance. The former seems necessary for any algorithm. The latter could be modeled as an out-of-distribution detector score, for which many methods are available (see e.g., [41]).
@@
Hedge [12] is a standard online learning algorithm which ensures sublinear regret when the number of hypotheses (in our case, the number of policies in $\Pi$) is finite. [^10] See Chapter 5 of [35] and Chapter 21 of [34] for modern introductions to Hedge. We would prefer not to assume that $\Pi$ is finite. Luckily, any policy $\Pi$ can be approximated within $\varepsilon$ when either (1) $\Pi$ has finite VC dimension and $\mathbf{s}$ $\sigma$-smooth or (2) $\Pi$ has finite Littlestone dimension. Thus we can run Hedge on this approximative policy class instead.
One other modification is necessary. In standard online learning, losses are observed on every time step, but our agent only receives feedback in response to queries. To handle this, we modify Hedge to only perform updates on time steps with queries and to issue a query with probability $p$ on each time step. Continuing our lucky streak, [33] analyzes exactly this modification of Hedge.
@@
10:     for $t$ from $1$ to $T$ do $\triangleright$ Run one step of Hedge, which selects policy $\pi_{t}$
@@
13:              Query mentor and observe $\pi^{m}(s_{t})$
14:              $\ell(t,\pi)\leftarrow\mathbf{1}(\pi(s_{t})\neq\pi^{m}(s_{t}))$ for all $\pi\in\tilde{\Pi}$
@@
21:         if $\min_{(s,a)\in X:a=\pi_{t}(s_{t})}||s_{t}-s||>\varepsilon^{1/n}$ then $\triangleright$ Ask for help if out-of-distribution
22:              Query mentor and observe $\pi^{m}(s_{t})$
23:              $X\leftarrow X\cup\{(s_{t},\pi^{m}(s_{t}))\}$
@@
Let $A=\{0,1\}$. Assume $\pi^{m}\in\Pi$ where either (1) $\Pi$ has finite VC dimension $d$, $\mathbf{s}$ is $\sigma$-smooth, and $\varepsilon T\log T>12\sigma d\log(4e^{2}/\varepsilon)$ or (2) $\Pi$ has finite Littlestone dimension $d$. Then for any $T\in\mathbb{N}$ and $\varepsilon>0$, Algorithm 1 satisfies
@@
$\operatorname*{\mathbb{E}}[|Q_{T}|]\in$ $\ O\left(\sqrt{\frac{T}{\varepsilon}}+\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T+\frac{\operatorname*{diam}(\mathbf{s})^{n}}{\varepsilon}\right)$
In Case 1, the expectation is over the randomness of both $\mathbf{s}$ and the algorithm, while in Case 2, the expectation is over only the randomness of the algorithm. Also, $R_{T}$ and $Q_{T}$ clearly have no dependence on $\sigma$ in Case 2, but we include $\sigma$ anyway to avoid writing two separate bounds.
@@
Let $A=\{0,1\}$. Assume $\pi^{m}\in\Pi$ where either (1) $\Pi$ has finite VC dimension $d$ and $\mathbf{s}$ is $\sigma$-smooth or (2) $\Pi$ has finite Littlestone dimension $d$. Then for any $T\in\mathbb{N}$, Algorithm 1 with $\varepsilon=T^{\frac{-2n}{2n+1}}$ satisfies
@@
$\operatorname*{\mathbb{E}}[|Q_{T}|]\in$ $\ O\left(T^{\frac{4n+1}{4n+2}}\left(\frac{d}{\sigma}\log T+\operatorname*{diam}(\mathbf{s})^{n}\right)\right)$
We can also make the query bound unconditional in exchange for slightly worse bounds. The probability that $|Q_{T}|$ asymptotically exceeds the bound from Theorem 5.2 goes to 0, so the query threshold in Corollary 5.2.1 is almost never reached. Thus we still obtain subconstant regret.
Corollary 5.2.1.
If Algorithm 1 is modified to stop querying after $T^{\frac{4n+1.5}{4n+2}}(\frac{d}{\sigma}\log T+\operatorname*{diam}(\mathbf{s})^{n})$ queries, then under the conditions of Theorem 5.2, we have $\operatorname*{\mathbb{E}}[R_{T}]\in O\big{(}\frac{dL}{\sigma}T^{\frac{-1}{8n+4}}\log T\big{)}$.

### v4

@@ 3 Model
If $\mu_{t}(x_{t},y_{t})\in[0,1]$ is the chance of avoiding catastrophe on time step $t$ (conditioned on no prior catastrophe), then $\prod_{t=1}^{T}\mu_{t}(x_{t},y_{t})$ is the agent’s overall chance of avoiding catastrophe by the chain rule of probability. For given $\boldsymbol{x},\boldsymbol{y},\boldsymbol{\mu},\pi^{m}$, the agent’s regret is
$R_{T}(\boldsymbol{x},\boldsymbol{y},\boldsymbol{\mu},\pi^{m})=\prod_{t=1}^{T}\mu_{t}^{m}(x_{t})-\prod_{t=1}^{T}\mu_{t}(x_{t},y_{t})$
We write $R_{T}=R_{T}(\boldsymbol{x},\boldsymbol{y},\boldsymbol{\mu},\pi^{m})$ for brevity. We desire subconstant worst-case regret: the total (not average) expected regret should go to 0 for any $\boldsymbol{\mu}$ and $\pi^{m}$; formally, $\lim_{T\to\infty}\sup_{\boldsymbol{\mu},\pi^{m}}\operatorname*{\mathbb{E}}[R_{T}]=0$.
The interpretation of a bound on $\operatorname*{\mathbb{E}}[R_{T}]$ depends on the quality of the mentor. For example, subconstant regret becomes trivial if $\lim_{T\to\infty}\operatorname*{\mathbb{E}}\big{[}\prod_{t=1}^{T}\mu_{t}^{m}(x_{t})\big{]}=0$. However, we think that high-stakes AI applications should ensure the presence of a mentor who is almost always safe, i.e., $\operatorname*{\mathbb{E}}\big{[}\prod_{t=1}^{T}\mu_{t}^{m}(x_{t})\big{]}\approx 1$. If no such mentor exists for some application, perhaps it is better to avoid the application altogether. Also, our regret bounds include rates of convergence, so even if the mentor policy will eventually cause catastrophe, we can still bound how quickly the agent becomes unsafe.
@@
VC dimension (Vapnik & Chervonenkis, 1971) and Littlestone dimension (Littlestone, 1988) are standard measures of learning difficulty which capture the ability of a hypothesis class (in our case, a policy class) to realize arbitrary combinations of labels (in our case, actions). We omit the precise dimensions since we only utilize these concepts via existing results. See Shalev-Shwartz & Ben-David (2014) for a comprehensive overview.
@@
The diameter of a set $X\subseteq\mathcal{X}$ is defined by $\operatorname*{diam}(X)=\max_{x,x^{\prime}\in X}||x-x^{\prime}||$. All logarithms and exponents are base $e$ unless otherwise noted.
@@ 4 Avoiding catastrophe is impossible in general
We begin by showing that in general, any algorithm with sublinear mentor queries has arbitrary poor regret in the worst-case, even when inputs are i.i.d. on $[0,1]$. The result also holds even if the algorithm knows $L$ and $\boldsymbol{x}$ ahead of time. The proof can be found in Appendix A.
@@
The worst-case expected regret of any algorithm with sublinear queries goes to 1 as $T$ goes to infinity. Formally, $\lim_{T\to\infty}\ \,\sup_{\boldsymbol{\mu},\pi^{m}}\,\operatorname*{\mathbb{E}}[R_{T}]=1$.
@@ 4.1 Intuition
We partition $\mathcal{X}$ into equally-sized sections that are “independent” in the sense that querying an input in section $i$ gives you no information about section $j$. The number of sections is determined by a function $f:\mathbb{N}\to\mathbb{N}$ that we will choose. If $|Q_{T}|\in o(f(T))$, most of these sections will never contain a query. When the agent sees an input in a section not containing a query, it essentially has to guess, meaning it will be wrong a constant fraction of the time. We then choose a payoff function (which is the same for all time steps) which makes the wrong guesses as costly as possible, subject to the local generalization constraint. Figure 2 fleshes out this idea.
@@
A natural idea is to try $f(T)=T$, but this doesn’t quite work: even if the agent chooses wrong on every time step, the minimum payoff is still at least $1-\frac{L}{2T}$, and $\lim_{T\to\infty}\prod_{t=1}^{T}\left(1-\frac{L}{2T}\right)=\lim_{T\to\infty}\left(1-\frac{L}{2T}\right)^{T}=e^{-L/2}$. In order for the regret to approach 1, we need $f(T)$ to be asymptotically between $|Q_{T}|$ and $T$ (such $f$ must exist since $|Q_{T}|\leq g(T)\in o(T)$). This leads to the following bound: $\prod_{t=1}^{T}\mu(x_{t},y_{t})\leq\big{(}1-\frac{L}{\Theta(f(T))}\big{)}^{\Theta(T)}$. When $f(T)\in o(T)$, the right hand side converges to 0, while $\prod_{t=1}^{T}\mu^{m}(x_{t})=1$. In words, the agent is nearly guaranteed to cause catastrophe, despite the existence of a policy which is guaranteed to avoid catastrophe.
$x$$\mu$  $\mu(x,0)$$\mu(x,1)$$\cfrac{L}{2f(T)}$$\frac{1}{f(T)}$$\frac{2}{f(T)}$$\frac{3}{f(T)}$$\frac{f(T)-1}{f(T)}$$\dots$1001111
Figure 2: An illustration of the construction we use to prove Theorem 4.1 (not to scale). The horizontal axis indicates the input $x\in[0,1]$ and the vertical axis indicates the payoff $\mu(x,y)\in[0,1]$. The solid line represents $\mu(x,0)$ and the dotted line represents $\mu(x,1)$. In each section, one of the actions has the optimal payoff of 1, and the other action has the worst possible payoff allowed by $L$, reaching a minimum of $1-\frac{L}{2f(T)}$. Crucially, both actions result in a payoff of 1 at the boundaries between sections: this allows us to “reset” for the next section. As a result, we can freely toggle the optimal action for each section independently.
@@ 4.2 Formal definition of construction
Let $\mathcal{X}=[0,1]$ and $\mathcal{D}_{t}=U$ for each $t\in[T]$. Assume that $L\leq 1$; this will simplify the math and only makes the problem easier for the agent. We define a family of payoff functions parameterized by a function $f:\mathbb{N}\to\mathbb{N}$ and a bit string $\boldsymbol{a}=(a_{1},a_{2},\dots,a_{f(T)})\in\{0,1\}^{f(T)}$. The bit $a_{j}$ will denote the optimal action in section $j$. Note that $f(T)\geq 1$ and since we defined $\mathbb{N}$ to exclude 0.
@@
$\mu_{f,\boldsymbol{a}}(x,y)=\begin{cases}1&\textnormal{ if }y=a_{j(x)}\\ 1-L\left(\cfrac{1}{2f(T)}-|m_{j(x)}-x|\right)&\textnormal{ if }y\neq a_{j(x)}\\ \end{cases}$
We will use this payoff function for all time steps: $\mu_{t}=\mu_{f,\boldsymbol{a}}$ for all $t\in[T]$. Let $\pi^{m}$ be any policy which is optimal for $\mu_{f,\boldsymbol{a}}$. Note that there is a unique optimal action for each $x_{t}$, since each $x_{t}$ belongs to exactly one $X_{j}$; formally, $\pi^{m}(x_{t})=a_{j(x_{t})}$.
For any $\boldsymbol{a}\in\{0,1\}^{f(T)}$, $\mu_{f,\boldsymbol{a}}$ is piecewise linear (trivially) and continuous (because both actions have payoff 1 on the boundary between sections). Since the slope of each piece is in $\{-L,0,L\}$, $\mu_{f,\boldsymbol{a}}$ is Lipschitz continuous. Thus by Proposition E.1, $\pi^{m}$ satisfies local generalization.
@@ 5 Avoiding catastrophe given finite VC or Littlestone dimension
Theorem 4.1 shows that avoiding catastrophe is impossible in general, which is also true in online learning. What if we restrict ourselves to settings where standard online learning is possible? Specifically, we assume that $\pi^{m}$ belongs to a policy class $\Pi$ where either (1) $\Pi$ has finite VC dimension $d$ and $\boldsymbol{x}$ is $\sigma$-smooth or (2) $\Pi$ has finite Littlestone dimension $d$ [^10] Recall from Section 1.3 that standard online learning becomes tractable under either of these assumptions.
This section presents a simple algorithm which guarantees subconstant regret and sublinear queries under either of those assumptions. Our algorithm needs to know $\Pi$, as is standard in online learning. The algorithm does not need to know $\sigma$ (in the smooth case) or $L$, and can handle an unbounded input space (the number of queries simply scales with the maximum distance between observed inputs).
For simplicity, we initially prove our result for $\mathcal{Y}=\{0,1\}$ (the proof can be found in Appendix B). Appendix C extends our result to many actions using the standard “one versus rest” reduction. [^11] For each action $y$, we learn a binary classifier which predicts whether $\pi^{m}(x)=y$. If every binary classifier is correct, we can correctly determine $\pi^{m}(x)$. See, e.g., Chapter 29 of Shalev-Shwartz & Ben-David (2014).
@@ 5.1 Intuition behind the algorithm
Algorithm 1 has two simple components: (1) run a modified version of the Hedge algorithm for online learning, but (2) ask for help for unfamiliar inputs (specifically, when the input is very different from any queried input with the same action under the proposed policy). Hedge ensures that the number of time steps where the agent’s action doesn’t match the mentor’s is small, and asking for help for unfamiliar inputs ensures that when we do make a mistake, the cost isn’t too high. This algorithmic structure seems quite natural: mostly follow a baseline strategy, but ask for help when out-of-distribution.
Agnostic to feature embedding.
The algorithm does not need to know how inputs are represented in $\mathbb{R}^{n}$, since the only nontrivial operation it performs on the input space is computing a nearest neighbor distance (line 21 of Algorithm 1). This could be modeled as an out-of-distribution detector score, for which many methods are available (see e.g., Yang et al. (2024)). This agnosticism reduces the burden on the local generalization assumption: it suffices that there exists some feature embedding satisfying this assumption.
@@
Hedge (Freund & Schapire, 1997) is a standard online learning algorithm which ensures sublinear regret when the number of hypotheses (in our case, the number of policies in $\Pi$) is finite. [^12] See Chapter 5 of Slivkins et al. (2019) and Chapter 21 of Shalev-Shwartz & Ben-David (2014) for modern introductions to Hedge. We would prefer not to assume that $\Pi$ is finite. Luckily, any policy $\Pi$ can be approximated within $\varepsilon$ when either (1) $\Pi$ has finite VC dimension and $\boldsymbol{x}$ is $\sigma$-smooth or (2) $\Pi$ has finite Littlestone dimension. Thus we can run Hedge on this approximative policy class instead.
One other modification is necessary. In standard online learning, losses are observed on every time step, but our agent only receives feedback in response to queries. To handle this, we modify Hedge to only perform updates on time steps with queries and to issue a query with probability $p$ on each time step. Continuing our lucky streak, Russo et al. (2024) analyze exactly this modification of Hedge.
@@
10:     for $t$
from
$1$
to
$T$ do $\triangleright$ Run one step of Hedge, which selects policy $\pi_{t}$
@@
13:              Query mentor and observe $\pi^{m}(x_{t})$
14:              $\ell(t,\pi)\leftarrow\mathbf{1}(\pi(x_{t})\neq\pi^{m}(x_{t}))$ for all $\pi\in\tilde{\Pi}$
@@
21:         if $\min_{(x,y)\in X:y=\pi_{t}(x_{t})}||x_{t}-x||>\varepsilon^{1/n}$ then $\triangleright$ Ask for help if out-of-distribution
22:              Query mentor and observe $\pi^{m}(x_{t})$
23:              $X\leftarrow X\cup\{(x_{t},\pi^{m}(x_{t}))\}$
@@
Let $\mathcal{Y}=\{0,1\}$. Assume $\pi^{m}\in\Pi$ where either (1) $\Pi$ has finite VC dimension $d$, $\boldsymbol{x}$ is $\sigma$-smooth, and $\varepsilon T\log T>12\sigma d\log(4e^{2}/\varepsilon)$ or (2) $\Pi$ has finite Littlestone dimension $d$. Then for any $T\in\mathbb{N}$ and $\varepsilon>0$, Algorithm 1 satisfies
@@
$\operatorname*{\mathbb{E}}[|Q_{T}|]\in$ $\ O\left(\sqrt{\frac{T}{\varepsilon}}+\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T+\frac{\operatorname*{diam}(\boldsymbol{x})^{n}}{\varepsilon}\right)$
In Case 1, the expectation is over the randomness of both $\boldsymbol{x}$ and the algorithm, while in Case 2, the expectation is over only the randomness of the algorithm. Also, $R_{T}$ and $Q_{T}$ clearly have no dependence on $\sigma$ in Case 2, but we include $\sigma$ anyway to avoid writing two separate bounds.
@@
Let $\mathcal{Y}=\{0,1\}$. Assume $\pi^{m}\in\Pi$ where either (1) $\Pi$ has finite VC dimension $d$ and $\boldsymbol{x}$ is $\sigma$-smooth or (2) $\Pi$ has finite Littlestone dimension $d$. Then for any $T\in\mathbb{N}$, Algorithm 1 with $\varepsilon=T^{\frac{-2n}{2n+1}}$ satisfies
@@
$\operatorname*{\mathbb{E}}[|Q_{T}|]\in$ $\ O\left(T^{\frac{4n+1}{4n+2}}\left(\frac{d}{\sigma}\log T+\operatorname*{diam}(\boldsymbol{x})^{n}\right)\right)$
Although our focus is the product of payoffs, Algorithm 1 also guarantees subconstant additive regret:
Theorem 5.3.
Let $\mathcal{Y}=\{0,1\}$. Assume $\pi^{m}\in\Pi$ where either (1) $\Pi$ has finite VC dimension $d$ and $\boldsymbol{x}$ is $\sigma$-smooth or (2) $\Pi$ has finite Littlestone dimension $d$. Then for any $T\in\mathbb{N}$, Algorithm 1 with $\varepsilon=T^{\frac{-2n}{2n+1}}$ satisfies
$\operatorname*{\mathbb{E}}\left[\sum_{t=1}^{T}\mu_{t}^{m}(x_{t})-\sum_{t=1}^{T}\mu_{t}(x_{t},y_{t})\right]\in O\left(\frac{dL}{\sigma}T^{\frac{-1}{2n+1}}\log T\right)$